OpenStudy (princeevee):

Newton's Second Law of Motion is expressed as F = ma. A lawn mower with constant mass is pushed over a length of grass. How much force is required to triple the acceleration rate? A. The force required is one-third the original force. B. The force required is three times the original force. C. The force required remains constant. D. The force required is four times the original.

OpenStudy (anonymous):

From F=ma you can get a=F/m so if you want 3 times a : \[3a=\frac{3F}{m}\] you need to triple the force!

OpenStudy (princeevee):

so the answer is B?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

F=ma a=F/m 3a = 3F/m Triple force is required to triple the acceleration